Aggregator

MOLLER install 11/5/2026

6 hours 53 minutes ago
Log Number: 4551860Logbook(s): Tags: Entry Makers: jones@jlab.orgBody: 

o Cable tray for Spectrometer IC cables installed. IC rack wire is complete.
o Cleaned up hall for tour
o Transported the TM0 Water cooled leads to testlab
o Welding of one cable tray support stand done.

jones

VWien + HWien positive and negative plates set at +/- 19 kV for overnight soak

8 hours 6 minutes ago
Log Number: 4551856Logbook(s): Entry Makers: mamunBody: 

At the end of ~4 h soak with VWie+HWien negative supplies at -19 kV, positive supplies of VWien + HWien are ramped to +19 kV in 1 kV steps every 30 sec. No vacuum activity is observed during the ramping nor during the last 4 h soak with negative plates at -19 kV. Tomorow morning we will ramp down both Wiens HV and turn off for Wien magnet yoke removal work.

mamun

Hall C daily summary

8 hours 43 minutes ago
Log Number: 4551852Logbook(s): Body: 

Today Hall C technicians worked on.

Fabricating NUCAT lead shielding frame assembly
Machining MOLER Barite shield wall assembly parts.
Printing and test-fitting and modifying NUCAT parts
Assembling filament extrusion project.
Fabricating poly shielding assembly for NUCAT
No cross division work planned for 11/06/26

jnines

Cryogenics Work Update

9 hours 9 minutes ago
Log Number: 4551850Logbook(s): Body: 

• The CHL2’s 2K Cold Box lid installed, late afternoon started initial vacuum pumping
• HDR mounting hardware and actuator for converting MV381 and MV368 to conversion control valves- in progress
• CHL Control Room new monitors wall mount installation - in progress
• Relocated old CHL Control Room 19" racks to building 36 (qty-4)
• SL PLC 6/8 pin cable harness - in progress
• Staged new spare MECOS drive for CHL1’s 2K Cold Box in preparation for Power up/testing - in progress

jwilson

iocsoftnl08rf restored

9 hours 46 minutes ago
Log Number: 4551844Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-11-04 12:10:53 The output from the myRestore execution was as follows: iocsoftnl08rf <NOTE> rf3int: Provided macro *IOC* is never used <NOTE> rf3int: Provided macro *REBOOTPATH* is never used <NOTE> rf3hpa: Provided macro *IOC* is never used <NOTE> rf3hpa: Provided macro *REBOOTPATH* is never used <NOTE> rf3stp: Provided macro *IOC* is never used <NOTE> rf3stp: Provided macro *REBOOTPATH* is never used <NOTE> rf3fcc1: Provided macro *IOC* is never used <NOTE> rf3f
geesaman

iocsoftnl08rf restored

9 hours 47 minutes ago
Log Number: 4551843Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-11-03 15:10:53 The output from the myRestore execution was as follows: iocsoftnl08rf <NOTE> rf3int: Provided macro *IOC* is never used <NOTE> rf3int: Provided macro *REBOOTPATH* is never used <NOTE> rf3hpa: Provided macro *IOC* is never used <NOTE> rf3hpa: Provided macro *REBOOTPATH* is never used <NOTE> rf3stp: Provided macro *IOC* is never used <NOTE> rf3stp: Provided macro *REBOOTPATH* is never used <NOTE> rf3fcc1: Provided macro *IOC* is never used <NOTE> rf3f
geesaman

APPROVED Wien magnet work-CEBAF/Injector

10 hours 11 minutes ago
Log Number: 4551840Logbook(s): Body: 

Added this ATLis as related task to ATLis 121912 -- Coordination of CEBAF Wien filter Jobs.

ATLis 121914 -- Disconnect Injector Wien Magnets MWF1I04 and MWF1I06 -- linked to this one, and was declared complete a little over an hour ago, covering the DC Power disconnects.

PTW-12243 on-issue.

Task approved, added to SAM Cal beginning 11/9 for two weeks.

moser

COMMENT ON 26/27 SAM Power Supply Flushing

10 hours 19 minutes ago
Log Number: 4551839Logbook(s): Body: 

MSSG relayed that they completed NLSB BoxPS flushing (Dog1E) and will proceed with flushing the two supplies (TSB5C, TAG5C) in the the Tagger ServBldg (B200) tomorrow, Friday 11/6.

For the work in Tagger SB, the 480VAC outlet is needed. Coordinated with F&O and Hall D -- outlet was in-use to power a temporary VFD for a pump for the Hall D Chiller system. There are two pumps, and given the current maintenance status of Hall D they confirmed no issue operating on one pump until flushing wraps up tomorrow.

The 480VAC outlet is ready for MSSG as of now.

moser

COMMENT ON Upgrade 1L08 to a C75 zone.

10 hours 22 minutes ago
Log Number: 4551838Logbook(s): Body: 

Wired RF cables to the cross connect and made new, longer hoses for circulator LCW return. In the process of installing new LCW manifold fittings, some hard sediment buildup was noted on the removed fittings. Since things have been going at a good pace, we're taking the time to have the fittings that are staying in the machine cleaned to remove some of the gunk. We're aiming to have the water back on and leak checked by next week.

alexs

iocr999rffcc1 restored

11 hours 28 minutes ago
Log Number: 4551833Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-11-05 13:26:21 The output from the myRestore execution was as follows: iocr999rffcc1 <NOTE> Rf12Fcc_TestStand: Provided macro *IOC* is never used Rf12Fcc_TestStand c100_rffcc@2026-11-05 13:26 71(71) restored c100_ecal@2026-11-05 13:26 20(20) restored Execute: myRestore -execute -P c100_ecal -t R991_ECAL -m BLDGNUM=9 ZONE=9 CAV=1 (success) c100_ecal@2026-07-10 15:43 20(20) restored Execute: burtwb -r 3 -f /cs/opshome/burt/REBOOT/Rf12Fcc/snap-kc/R991-kc.snap (success) c100_
alarms

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Richardson
MCC Operators
Program Deputy Merz
Announcements

Access Status

Recent PSS History

recent pss history graph